1-Question 28 Convert the radian measure to degrees. Give answer using decimal degrees to the nearest hundredth. Use 3.1416 for

2-Question 30 Find all specined roots. Cube roots of 1. -1.1
Answer #1

1) we know that 3.1416 radian = 180°

So 2 radian = (2/3.1416)*180° = 114.59°

2)

Let, x^3 = 1

Then

x^3 - 1 = 0

Or, (x-1)(x^2+x+1) = 0

So, x -1 = 0 or, x^2 + x +1 = 0

So, x = 1 , -1/2 + √3i/2 , -1/2 - √3i/2

These are the cube root of 1